# Real-Time Dynamics in Two Dimensions with Tensor Network States via Time-Dependent Variational Monte Carlo

### v2

@@ Abstract
Reliably simulating two-dimensional many-body quantum dynamics with projected entangled pair states (PEPS) has long been a difficult challenge. In this work, we overcome this barrier for low-energy quantum dynamics by developing a stable and efficient time-dependent variational Monte Carlo (tVMC) framework for PEPS. By analytically removing all gauge redundancies of the PEPS manifold and exploiting tensor locality, we obtain a numerically well-conditioned stochastic reconfiguration equation amenable to robust solution using the efficient Cholesky decomposition, enabling long-time evolution in previously inaccessible regimes. We explain how the difficulties associated with the traditional PEPS algorithms for dynamics are naturally resolved within tVMC. We demonstrate the power and generality of the method through four representative real-time problems in two dimensions: (I) chiral edge propagation in a free-fermion Chern insulator; (II) fractionalized charge transport in a fractional Chern insulator; (III) vison confinement dynamics in the Higgs phase of a $\mathbb{Z}_{2}$ lattice gauge theory coupled to a Higgs field; and (IV) superfluidity and critical velocity in interacting bosons. All simulations are performed on $12\times 12$ or $13\times 13$ lattices with evolution times $T=10$ to $12$ using modest computational resources — 1 to 5 days on a single GPU card. Where exact benchmarks exist (case I), PEPS–tVMC matches free-fermion dynamics with high accuracy up to $T=12$. These results establish PEPS–tVMC as a practical and versatile tool for real-time quantum dynamics in two dimensions. The method significantly extends the reach of classical tensor-network simulations for studying elementary excitations in quantum many-body systems and provides a valuable computational counterpart to emerging quantum simulators.
@@ I Introduction
Understanding real-time dynamics of interacting quantum many-body systems in two dimensions (2D) is a fundamental challenge that spans condensed-matter physics, particle physics, and quantum simulation. While powerful classical methods exist for ground states—including tensor networks [1, 2] and quantum Monte Carlo [3, 4], real-time evolution in 2D remains notoriously difficult. Exact diagonalization is limited to small clusters; quantum Monte Carlo suffers from a dynamical sign problem. Recently, the combination of time-dependent variational Monte Carlo (tVMC) [5, 6, 7, 8, 9, 10] with neural quantum states (NQS) has emerged as a promising alternative [11, 12, 13, 14, 15, 16, 17]. NQS provide remarkable variational flexibility, while tVMC offers a principled formulation of real-time evolution. At present, extending NQS–tVMC to large 2D systems remains an active area of research, where one must balance expressive parametrizations with efficiency and stable optimization [18].
@@
Projected entangled pair states (PEPS) [19] provide a complementary variational family whose locality and multi-linearity make them particularly appealing for tVMC. Within VMC, only a single layer of PEPS tensors needs to be contracted, making the method substantially more efficient than conventional PEPS update schemes [20, 21, 22]. Furthermore, VMC allows parallel accumulation of samples, making the approach well suited for modern GPU architectures. In fact, PEPS-VMC has recently achieved considerable success for ground states, giving state-of-the-art results for the ground states of frustrated spins [23, 24, 25], fermion Hubbard model [26], and lattice gauge theory [27], owing to a foundational breakthrough in the sampling method of PEPS [28].
@@
Although PEPS obey an area law for entanglement entropy [29] and cannot represent volume-law states produced by fully thermalizing dynamics, they are well suited for local quenches in which the injected energy does not scale with system size. In this regime, the energy density remains zero (in the thermodynamic limit) and we expect the dynamics to preserve area-law entanglement [30]. Thus PEPS naturally targets low-energy local-quench dynamics — precisely the regime relevant for elementary excitations and real-time response functions, and directly related to experiments including scanning tunneling microscopy [31], angle resolved photoemission spectroscopy [32], particle scattering processes, and more. A controlled method capable of simulating such dynamics provides a valuable tool for studying 2D quantum systems.
@@
While dynamical simulations with matrix product states (MPS) are enormously successful in 1D [33, 34, 35, 36], the situation in 2D is far more challenging. In contrast to the extensive 1D literature, only a small number of works—represented by Refs. [37, 38, 39, 40, 41, 42, 43, 44, 45, 46], have reported real-time simulations using PEPS or iPEPS, illustrating the general difficulty of the problem. Traditional approaches to dynamical PEPS have relied on full-update-like methods based on double-layer contraction [20, 21, 22], which face several challenges: (a) high cost of double layer contraction, (b) difficulty in fixing the PEPS gauge freedom, and (c) loss of positivity in the norm matrix approximately obtained from the contraction of the double-layer environment. We explain, in Sec. III.4, that the last two issues are naturally resolved while the first one is significantly alleviated within PEPS–tVMC, at the expense of manageable statistical noise.
@@
Despite the long development of PEPS [1], their combination with tVMC for real-time dynamics has received little attention, aside from our recent exploratory attempt in a simple setting [27]. Here, we develop PEPS–tVMC systematically, analyze its structure, and demonstrate its ability to simulate large-scale 2D dynamics. Our approach builds on stochastic reconfiguration (SR) [5], identifies and removes the gauge redundancy of PEPS parameters, and exploits tensor locality to simplify and precondition the SR equations. After gauge removal, the SR equation exhibits remarkable numerical stability: the SR residual remains as low as $10^{-9}$ to $10^{-25}$ in the examples studied. The central message of this work is that by removing PEPS gauge freedom and exploiting locality, the tVMC equations become surprisingly stable to solve, and that, as a result, PEPS—long viewed as difficult for real-time evolution—can in fact provide a powerful and controllable variational framework for 2D dynamics. This resolves long-standing challenges that have hindered dynamical simulations with PEPS since their introduction [19].
@@ II Time-Dependent Variational Monte Carlo
Consider a many-body lattice quantum system with computational basis $\bm{\mathbf{s}}=(s_{1},\dots,s_{N})$. To perform a variational simulation, we adopt a holomorphic parametrization of the unnormalized quantum state $\Psi_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})$ with $N_{p}$ complex variational parameters $\bm{\mathbf{\theta}}=\{\theta_{\alpha}:\alpha=1,2,\dots,N_{p}\}$. At time $t$, the exact evolution under a Hamiltonian $H$ for a short interval $dt$ induces an evolution from $\bm{\mathbf{\theta}}(t)$ to $\bm{\mathbf{\theta}}(t+dt)$ that minimizes the Fubini–Study distance [47] between $\Psi_{\bm{\mathbf{\theta}}(t+dt)}$ and $e^{-iHdt}\Psi_{\bm{\mathbf{\theta}}(t)}$. This yields the SR equation:
@@ III Stochastic reconfiguration with PEPS
A PEPS defines a variational wavefunction $\Psi_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})$ parametrized by the tensor elements $A[\bm{\mathbf{x}}]^{p}_{lrdu}$ at each lattice site $\bm{\mathbf{x}}$, as illustrated in Fig. 1(a,b). Each index tuple $(\bm{\mathbf{x}},p,l,r,d,u)$ corresponds to one variational parameter, replacing the abstract index $\alpha$ introduced earlier. The physical index $p$ labels the local Hilbert-space configuration (e.g., $p=0,1$ for vacuum or occupation), while the virtual indices $(l,r,d,u)$ take values from $1$ to $D$ (except on boundaries).
@@
Hence every gauge vector also lies in the null space of $S=\mathcal{O}^{\dagger}\mathcal{O}$, and must be removed before solving the SR equation
@@
The $2L(L-1)D^{2}$ gauge vectors are not linearly independent. Consider a closed loop of virtual bonds and take $E^{ij}$ to be the identity matrix on each bond. The sum of the corresponding gauge vectors vanishes, indicating linear dependence. These relations are generated by the $(L-1)^{2}$ elementary plaquettes of the square lattice. We denote the associated constraint vectors by $\bm{\mathbf{w}}$, which are known exactly in the basis of the gauge vectors and independent of tensor content.
@@ III.1 Gauge-fixing within SR
We now outline the procedure for eliminating null vectors within the SR formulation. First, we construct a projection matrix $P$ onto the subspace orthogonal to the constraint vectors $\bm{\mathbf{w}}$, thereby reducing the number of independent gauge vectors to
@@
We then form an $N_{p}\times(N_{\rm gv}+2)$ matrix $T$ whose columns consist of these reduced gauge vectors together with $\bm{\mathbf{u}}_{1}$ and $\bm{\mathbf{u}}_{2}$. Performing a complete QR decomposition $T=\widetilde{Q}R$, we define $Q$ as the last $N_{p}-N_{\rm gv}-2$ columns of $\widetilde{Q}$. Since $T$ has full column rank, the remaining columns of $\widetilde{Q}$ span the orthogonal complement to the column space of $T$. By construction, $Q$ projects onto the subspace orthogonal to all null directions.
The SR equation is then solved in this reduced space:
@@ III.2 Gauge-fixing within minSR
More generally, the null vector of $\mathcal{O}^{\dagger}$ is the weight vector $W$ used in performing the statistical average: $\braket{A}\approx\sum_{s=1}^{N_{s}}W_{s}A_{s}/N_{s}$, which in our case is simply the identity weight. A non-trivial weight vector would arise if reweighing is used in the Monte Carlo estimate. The removal of this null vector may be also benificial to NQS-tVMC in the minSR form.
Finally, we comment on the choice between SR and minSR. As the system size gets larger, $N_{p}$ eventually overtakes $N_{s}$ and one should use minSR for efficiency. However, it is always advised to study a problem first at small size and ideally benchmark against exact methods, in which case typically $N_{s}>N_{p}$ and it is only correct to use SR.
@@ III.3 Small-$o$ trick in minSR
In tVMC, the memory bottleneck is storing $O$, which can be a serious issue on GPUs. For a PEPS with local dimension $d$, this “small-oo” trick reduces memory cost by a factor of $d$. For the $\mathbb{Z}_{N}$ gauge-invariant PEPS used in Sec. IV.3, the reduction is very substantial—a factor of $dN^{3}$. In practice, this saving often determines whether a large-scale GPU simulation is doable at all.
@@ III.4 Comparison with double-layer methods
We briefly discuss the difficulties of double-layer methods mentioned in the introduction: computational cost, gauge fixing, and positivity of the norm matrix. Although manifested differently, they reflect the same underlying difficulty.
@@
Normalization.
Because PEPS $\psi$ is an unnormalized ansatz, to perform a variational update, one needs to also form the norm matrix $N$ accounting for the norm $\braket{\psi|\psi}$. In practice, approximate contraction leads to violations of positivity of $N$, which destabilizes the generalized eigenvalue problems used in variational updates, especially in time evolution. Tiny errors in the boundary contraction can lead to large deviations in the time trajectory. tVMC avoids this issue entirely: normalization is enforced simply by subtracting $O_{\rm avg}$ and $h_{\rm avg}$ in Eq. 5, requiring no explicit norm matrix.
In the end, we mention that any conservation law in the computational basis can be implemented, in VMC, by sampling in the subsector of configurations with a definite total charge. This is equivalent to acting to the state a global projector of the desired charge sector, which significantly increases the representability of the underlying bare PEPS.
@@ IV Results
We now apply PEPS–tVMC to study four distinct two-dimensional models—bosonic and fermionic, free and interacting, and including systems with gauge symmetry or nontrivial topology. PEPS configurations are sampled using the sequential sampling algorithm, which was first introduced in Ref. [28], and later proved to satisfy detailed balance in Ref. [52]. The local energies and log-derivatives are computed via the contraction of the single-layer network shown in Fig. 1(d) (see Refs. [28, 52, 27] for methodological details). All simulation parameters are given in the Supplementary Material [50].
@@ IV.2 Bosonic fractional quantum Hall effect
We consider the $1/8$ density, where the lowest magnetic band is half filled, corresponding to filling $\nu=1/2$. With the filling belonging to the Jain sequence [57], composite fermion theory [58, 59] predicts that the ground state to be a bosonic analogue of the Laughlin state [56] with quasi-particle charge $1/2$. The lattice structure [60] also complicates the theory which is initially developed for the continuum. We now explicitly check the fractional transport in real-time.
To account for the finite boundary, in practice, we use particle number $(L_{x}-1)(L_{y}-1)/8\approx 15$  [^1] We thank Bartholomew Andrews for letting us know this practice.. As shown in Fig. 3 (a), the ground state VMC gives a charge distribution that is uniformly $1/8$ in the bulk.
@@
To probe the fractional charge transport, we thread an external time-dependent flux $\Phi(t)$ through the central plaquette, cf. the red $\otimes$ in Fig. 3 (a), increasing linearly from 0 to $2\pi$ over a period $T_{\rm flux}$. The external flux is implemented via a Peierls phase on the red horizontal bonds shown in Fig. 3 (a). This generates a steady electromotive force around any closed loop enclosing the plaquette and induces a Hall transport from the middle of the system to the edge. If the system is indeed a fractional quantum Hall state, a quantized charge of $\Delta Q=\nu=1/2$ is expected to be transported per pumping cycle [62, 63]. Indeed, Fig. 3 (b) shows that PEPS–tVMC confirms this quantized transfer of charge. A linear fit of the charges within a central region, cf. the blue circle in Fig. 3 (a), versus $t/T_{\text{flux}}$ gives a slope of $0.5053$, close to the expected Hall response of $1/2$.
@@
This result provides a direct real-time confirmation of the composite-fermion prediction of fractional quasiparticle charge. While previous numerical studies of fractional Chern insulators relied primarily on static diagnostics, such as entanglement spectra or density profiles of pinned quasiholes, PEPS–tVMC enables us to directly visualize the transport of fractional charge in two dimensions. This establishes a dynamical counterpart to the usual static signatures of topological order. It would be exciting to study other anyon dynamics with PEPS–tVMC in the future.

### v3

@@ Abstract
Reliably simulating two-dimensional many-body quantum dynamics with projected entangled pair states (PEPS) has long been a difficult challenge. In this work, we overcome this barrier for low-energy quantum dynamics by developing a stable and efficient time-dependent variational Monte Carlo (tVMC) framework for PEPS. By analytically removing all gauge redundancies of the PEPS manifold and exploiting tensor locality, we obtain a numerically well-conditioned stochastic reconfiguration equation amenable to robust solution using the efficient Cholesky decomposition, enabling long-time evolution in previously inaccessible regimes. We explain how the difficulties associated with the traditional PEPS algorithms for dynamics are naturally resolved within tVMC. We demonstrate the power and generality of the method through four representative real-time problems in two dimensions: (I) chiral edge propagation in a free-fermion Chern insulator; (II) fractionalized charge transport in a fractional Chern insulator; (III) vison confinement dynamics in the Higgs phase of a $\mathbb{Z}_{2}$ lattice gauge theory coupled to a Higgs field; and (IV) superfluidity and critical velocity in interacting bosons. All simulations are performed on $12\times 12$ or $13\times 13$ lattices with evolution times $T=10$ to $12$ using modest computational resources — 1 to 5 days on a single GPU card. Where exact benchmarks exist (case I), PEPS–tVMC matches free-fermion dynamics with high accuracy up to $T=12$. These results demonstrate PEPS–tVMC as a practical and versatile tool for real-time quantum dynamics in two dimensions. The method significantly extends the reach of classical tensor-network simulations for studying elementary excitations in quantum many-body systems in real-time and provides a valuable computational counterpart to emerging quantum simulators.
@@ I Introduction
Understanding real-time dynamics of interacting quantum many-body systems in two dimensions (2D) is a fundamental challenge that spans condensed-matter physics, particle physics, and quantum simulation. While powerful classical methods exist for ground states—including tensor networks [1, 2] and quantum Monte Carlo [3, 4], real-time evolution in 2D remains notoriously difficult. Exact diagonalization is limited to small clusters; quantum Monte Carlo suffers from a dynamical sign problem. Recently, the combination of time-dependent variational Monte Carlo (tVMC) [5, 6, 7], a continuous-time generalization of the stochastic reconfiguration method [8, 9, 10], with neural quantum states (NQS) [11] has emerged as a promising alternative [12, 13, 14, 15, 16, 17, 18]. NQS provide remarkable variational flexibility, while tVMC offers a principled formulation of real-time evolution. At present, extending NQS–tVMC to large 2D systems remains an active area of research, where one must balance expressive parametrizations with efficient and stable optimization [19].
@@
Projected entangled pair states (PEPS) [20] provide a complementary variational family whose locality and multi-linearity make them particularly appealing for tVMC. Within VMC, only a single layer of PEPS tensors needs to be contracted, making the method substantially more efficient than conventional PEPS update schemes [21, 22, 23]. Furthermore, VMC allows parallel accumulation of samples, making the approach well suited for modern GPU architectures. In fact, PEPS-VMC has recently achieved considerable success for ground states, giving competitive results for the ground states of frustrated spins [24, 25, 26], fermion Hubbard model [27], and lattice gauge theory [28], owing to a foundational breakthrough in the sampling method of PEPS [29].
@@
Although PEPS obey an area law for entanglement entropy [30] and cannot represent volume-law states produced by fully thermalizing dynamics, they are well suited for local quenches in which the injected energy does not scale with system size. In this regime, the energy density remains zero (in the thermodynamic limit) and we expect the dynamics to preserve area-law entanglement [31, 32]. Thus PEPS naturally targets low-energy local-quench dynamics — precisely the regime relevant for elementary excitations and real-time response functions, and directly related to experiments including scanning tunneling microscopy [33], angle resolved photoemission spectroscopy [34], particle scattering processes, and more. A controlled method capable of simulating such dynamics provides a valuable tool for studying 2D quantum systems.
@@
While dynamical simulations with matrix product states (MPS) are enormously successful in 1D [35, 36, 37, 38], the situation in 2D is far more challenging. In contrast to the extensive 1D literature, only a small number of works—represented by Refs. [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49], have reported real-time simulations using PEPS or iPEPS, illustrating the general difficulty of the problem. Traditional approaches to dynamical PEPS have relied on full-update-like methods based on double-layer contraction [21, 22, 23], which face several challenges: (a) high cost of double-layer contraction, (b) difficulty in fixing the PEPS gauge freedom, and (c) loss of positivity in the quantum metric approximately obtained from the contraction of the double-layer environment. In Sec. III.4, we explain that, within PEPS–tVMC, the last two issues are naturally resolved while the first one is significantly alleviated, at the expense of manageable statistical noise.
@@
Despite the long development of PEPS [1], their combination with tVMC for real-time dynamics has received virtually no attention, aside from our recent exploratory attempt in a simple setting [28]. Here, we develop PEPS–tVMC systematically, analyze its structure, and demonstrate its ability to simulate large-scale 2D dynamics. Our approach builds on stochastic reconfiguration (SR) [8], identifies and removes the gauge redundancy of PEPS parameters, and exploits tensor locality to simplify and precondition the SR equations. After gauge removal, the SR equation exhibits remarkable numerical stability: the SR residual remains as low as $10^{-9}$ to $10^{-25}$ in the examples studied. The central message of this work is that by removing PEPS gauge freedom and exploiting locality, the tVMC equations become remarkably stable to solve, and that, as a result, PEPS—long viewed as difficult for real-time evolution—can in fact provide a powerful and controllable variational framework for 2D dynamics. This resolves long-standing challenges that have hindered dynamical simulations with PEPS since their introduction [20].
@@ II Time-Dependent Variational Monte Carlo
Consider a many-body lattice quantum system with computational basis $\bm{\mathbf{s}}=(s_{1},\dots,s_{N})$. To perform a variational simulation, we adopt a holomorphic parametrization of the unnormalized quantum state $\Psi_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})$ with $N_{p}$ complex variational parameters $\bm{\mathbf{\theta}}=\{\theta_{\alpha}:\alpha=1,2,\dots,N_{p}\}$. At time $t$, the exact evolution under a Hamiltonian $H$ for a short interval $dt$ induces an evolution from $\bm{\mathbf{\theta}}(t)$ to $\bm{\mathbf{\theta}}(t+dt)$ that minimizes the Fubini–Study distance [50] between $\Psi_{\bm{\mathbf{\theta}}(t+dt)}$ and $e^{-iHdt}\Psi_{\bm{\mathbf{\theta}}(t)}$. Taking $dt$ to 0, this yields the SR equation for tVMC [5]:
@@ III Stochastic reconfiguration with PEPS
A PEPS defines a variational wavefunction $\Psi_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})$ parametrized by the tensor elements $A[\bm{\mathbf{x}}]^{p}_{lrdu}$ at each lattice site $\bm{\mathbf{x}}$, as illustrated in Fig. 1(a,b). Each index tuple $(\bm{\mathbf{x}},p,l,r,d,u)$ corresponds to one variational parameter, replacing the abstract index $\alpha$ introduced earlier. The physical index $p$ labels the local Hilbert-space configuration (e.g., $p=0,1$ for vacuum or occupation), while the virtual indices $(l,r,d,u)$ take values from $1$ to $D$ (except on boundaries). The PEPS in this paper all have open boundary condition.
@@
Hence every gauge vector also lies in the null space of $S=\mathcal{O}^{\dagger}\mathcal{O}$, and must be removed before solving the SR equation.
@@
The $2L(L-1)D^{2}$ gauge vectors are not linearly independent. Consider a closed loop of virtual bonds and replace $E^{ij}$ with the identity matrix on each bond. The sum of the corresponding gauge vectors vanishes, indicating linear dependence. These relations are generated by the $(L-1)^{2}$ elementary plaquettes of the square lattice. We denote the associated constraint vectors by $\bm{\mathbf{w}}$, which are known exactly in the basis of the gauge vectors and independent of tensor content.
@@ III.1 Gauge-fixing within SR
We now outline the procedure for eliminating null vectors within the SR formulation. First, we construct a projection matrix onto the subspace orthogonal to the constraint vectors $\bm{\mathbf{w}}$, thereby reducing the number of independent gauge vectors to
@@
We then form an $N_{p}\times(N_{\rm gv}+2)$ matrix $T$ whose columns consist of these independent gauge vectors together with $\bm{\mathbf{u}}_{1}$ and $\bm{\mathbf{u}}_{2}$. Performing a complete QR decomposition $T=\widetilde{Q}R$, we define $Q$ as the last $N_{p}-N_{\rm gv}-2$ columns of $\widetilde{Q}$. Since $T$ has full column rank, the remaining columns of $\widetilde{Q}$ span the orthogonal complement to the column space of $T$. By construction, $Q$ projects onto the subspace orthogonal to all null directions.
The SR equation is then solved for $y$ in the reduced space:
@@ III.2 Gauge-fixing within minSR
More generally, the null vector of $\mathcal{O}^{\dagger}$ is the weight vector $W$ used in performing the statistical average: $\braket{A}\approx\sum_{s=1}^{N_{s}}W_{s}A_{\text{loc}}(s)/N_{s}$, which in our case is simply the identity weight. A non-trivial weight vector would arise if reweighting is used in the Monte Carlo estimate. The removal of this null vector may be also beneficial to NQS-tVMC in the minSR form.
Finally, we comment on the choice between SR and minSR. As the system size increases, $N_{p}$ eventually overtakes $N_{s}$ and one should use minSR for efficiency. However, it is always advised to study a problem first at small size and ideally benchmark against exact methods, in which case typically $N_{s}>N_{p}-N_{\text{gv}}-2$ and it is only correct to use SR.
@@ III.3 Small-$o$ trick in minSR
In tVMC, the memory bottleneck is storing $O$, which can be a serious issue on GPUs. For a PEPS with local dimension $d$, this “small-oo” trick reduces memory cost by a factor of $d$. For the $\mathbb{Z}_{N}$ gauge-invariant PEPS used in Sec. IV.3, the reduction is very substantial—a factor of $dN^{3}$. In practice, this saving often determines whether a large-scale GPU simulation is feasible at all.
@@ III.4 Comparison with double-layer methods
We briefly discuss the difficulties of double-layer methods mentioned in the introduction: computational cost, gauge fixing, and positivity of the quantum metric. Although manifested differently, they reflect the same underlying difficulty.
@@
Positivity.
Because PEPS are not in an isometric form, to perform a variational update, one needs to also form the non-trivial quantum metric of the parameters, induced from the inner product of the Hilbert space. This quantum metric is also known as the norm matrix $N$ in Ref. [23]. In practice, approximate contraction leads to violations of positivity of $N$, which destabilizes the generalized eigenvalue problems used in variational updates, especially in time evolution. Tiny errors in the boundary contraction can lead to large deviations in the time trajectory. tVMC avoids this issue entirely: normalization is enforced simply by subtracting $O_{\rm avg}$ and $h_{\rm avg}$ in Eq. 5, and the $S$ matrix in Eq. 2, serving as the quantum metric, is manifestly positive semi-definite.
In the end, we mention another reason why tVMC may be easier than double-layer methods. Any conservation law in the computational basis can be implemented, in VMC, by sampling in the subsector of configurations with a definite total charge. This is equivalent to acting on the state with a global projector of the desired charge sector, which significantly increases the representability of the underlying bare PEPS.
@@ IV Results
We now apply PEPS–tVMC to study four distinct two-dimensional models—bosonic and fermionic, free and interacting, and including systems with gauge symmetry or nontrivial topology. PEPS configurations are sampled using the sequential sampling algorithm, which was first introduced in Ref. [29], and later proved to satisfy detailed balance in Ref. [55]. The local energies and log-derivatives are computed via the contraction of the single-layer network shown in Fig. 1(d) (see Refs. [29, 55, 28] for methodological details). For Hamiltonians with no explicit time-dependence, the energies are observed to be well-conserved, with relative drift of $2\times 10^{-4}$ to $2\times 10^{-3}$ for the entire trajectory. We show the energy evolution and the simulation parameters in the Supplementary Material [53].
@@ IV.2 Bosonic fractional quantum Hall effect
We consider the $1/8$ density, where the lowest magnetic band is half filled, corresponding to filling $\nu=1/2$. With the filling belonging to the Jain sequence [60], composite fermion theory [61, 62] predicts that the ground state is a bosonic analogue of the Laughlin state [59] with quasi-particle charge $1/2$. The lattice structure [63] also complicates the theory which was initially developed for the continuum. We now explicitly check the fractional transport in real-time.
To account for the finite boundary, in practice, we use particle number $(L_{x}-1)(L_{y}-1)/8\approx 15$  [^1] We thank Bartholomew Andrews for bringing this practice to our attention.. As shown in Fig. 3 (a), the ground state VMC gives a charge distribution that is uniformly $1/8$ in the bulk.
@@
To probe the fractional charge transport, we thread an external time-dependent flux $\Phi(t)$ through the central plaquette, cf. the red $\otimes$ in Fig. 3 (a), increasing linearly from 0 to $2\pi$ over a period $T_{\rm flux}=10$. The external flux is implemented via a Peierls phase on the red horizontal bonds shown in Fig. 3 (a). This generates a steady electromotive force around any closed loop enclosing the plaquette and induces a Hall transport from the middle of the system to the edge. If the system is indeed a fractional quantum Hall state, a quantized charge of $\Delta Q=\nu=1/2$ is expected to be transported per pumping cycle [65, 66]. Indeed, Fig. 3 (b) shows that PEPS–tVMC confirms this quantized transfer of charge. A linear fit of the charges within a central region, cf. the blue circle in Fig. 3 (a), versus $t/T_{\text{flux}}$ gives a slope of $0.5053$, close to the expected Hall response of $1/2$.
@@
This result provides a direct real-time confirmation of the composite-fermion prediction of fractional quasiparticle charge. While previous numerical studies of fractional Chern insulators relied primarily on static diagnostics, such as entanglement spectra or density profiles of pinned quasiholes, PEPS–tVMC enables us to directly visualize the transport of fractional charge in two dimensions. This establishes a dynamical counterpart to the usual static signatures of topological order. It would be exciting to study other anyonic dynamics with PEPS–tVMC in the future.